# LlamaPIE [Uncaptioned image]: Proactive In-Ear Conversation Assistants

### v1

@@ Abstract
We introduce LlamaPIE, the first real-time proactive assistant designed to enhance human conversations through discreet, concise guidance delivered via hearable devices. Unlike traditional language models that require explicit user invocation, this assistant operates in the background, anticipating user needs without interrupting conversations. We address several challenges, including determining when to respond, crafting concise responses that enhance conversations, leveraging knowledge of the user for context-aware assistance, and real-time, on-device processing. To achieve this, we construct a semi-synthetic dialogue dataset and propose a two-model pipeline: a small model that decides when to respond and a larger model that generates the response. We evaluate our approach on real-world datasets, demonstrating its effectiveness in providing helpful, unobtrusive assistance. User studies with our assistant, implemented on Apple Silicon M2 hardware, show a strong preference for the proactive assistant over both a baseline with no assistance and a reactive model, highlighting the potential of LlamaPIE  to enhance live conversations. Code, dataset available at github.com/chentuochao/LlamaPIE
@@
Tuochao Chen, Nicholas Batchelder, Alisa Liu, Noah Smith, Shyamnath Gollakota University of Washington
@@ 1 Introduction
User interaction with language models (LMs) has primarily followed a turn-based dialogue format, where users actively request responses from LM assistants. This approach requires users to shift their attention, often carefully phrasing prompts to obtain useful answers. Thus, usability is restricted to scenarios where users can pause their activities to engage with the conversational assistant.
Figure 1: LlamaPIE is a Proactive, In-Ear assistant that augments human-to-human communication by providing discreet guidance via hearable devices. Its responses are short, provided only when helpful, and leverage the wearer’s memory of past events. In the figure, LlamaPIE assists Yuan by whispering 1-3 words to her during the conversation only when needed by anticipating user needs, remaining silent most of the time.
@@
Such proactive companions may enhance human interactions in a variety of contexts, including negotiations Deng et al. (2024), interviews Naim et al. (2018), customer support Deng et al. (2024), and cross-cultural communication Andolina et al. (2018). They also hold significant potential for supporting neurodiverse individuals, such as those with autism or social anxiety, by assisting in the interpretation of social cues Liu et al. (2024). They can also serve as personal memory assistants for older adults experiencing progressive memory loss, providing contextually relevant information during conversations Bermejo et al. (2020).
Designing proactive assistants presents several challenges. Modern language models are typically designed to provide detailed responses, but in our case, assistance should be limited to just 1–3 words and offered only when it enhances the conversation without disrupting its flow. The assistant must also be context-aware, leveraging the wearer’s memory of prior activities and interactions. Achieving this enhancement to live conversations requires real-time streaming processing, while privacy considerations necessitate processing conversational data on edge devices rather than sending it to the cloud. Finally, the lack of large, annotated datasets for in situ human conversations and real-time, hearable-based assistance presents a significant challenge for model training.
@@
Our work highlights a vision for proactive LM assistance that centers human-to-human conversation, and demonstrates the potential for such interactions to serve users in ways that complement the currently widespread human-AI “chat” interface.
@@ 2.2 Synthetic Dataset Generation
We construct examples by first creating the user memory (§2.2.2) and then the dialogue (§2.2.3). To improve the diversity and realisticness of the data, we draw from real conversational contexts and user profile datasets when synthetically generating data. The assistant should serve two main functions: providing reminders, e.g., helping users recall secondary details of events like names and places, and social guidance, e.g., helping the user continue the natural flow of the conversation. We use claude-3-5-sonnet-20240620 with Anthropic API to generate data (see §C.3 for dataset generation prompts).
@@ 2.2.2 Memory Generation or Selection
For synthetic memory, we use Claude to generate a user profile by providing five randomly selected keywords from a predefined list of 100 (see §C.2). Given these keywords, the model constructs a user profile and corresponding events. For SODA-based memory, instead of keywords, the model uses the context preceding SODA dialogues to generate the user profile. For PerLTQA-based memory, we select a random profile from the PerLTQA dataset, and two random events that are attached to that profile.
@@ 2.2.3 Dialogue Generation
The next step is dialogue generation, grounded in the user memory. We instruct Claude to construct a scenario that the user might encounter given their background. Each scenario is designed to fit a specified scenario type and illustrate a predefined use case, while also exemplifying two randomly selected principles of proactive assistants. To account for scenarios where users ignore assistance, some dialogues include an explicit instruction for Claude to generate interactions in which the user occasionally ignores the assistant’s messages (§C.3).
@@
Formatting for streaming and timing. To enable tokenized streaming and preserve timing information in dialogues — details typically absent in natural language data used for training LMs — we reformat Claude-generated dialogues to accommodate real-time processing. Gaps between speakers are replaced with silence markers, |SILENCE >, with each token representing 0.5 seconds of silence. Generated hesitation markers are replaced with silence markers as well (the hesitation duration is rounded up to the nearest equivalent number of tokens). This ensures that pauses and hesitations are accurately represented in the dataset, even when Claude’s generation does not include sufficiently long hesitations. Furthermore, since all gaps and hesitations are converting into the same silence marker, the streaming models do not have explicit marked information about hesitations. Future work could focus on more sophisticated modeling of conversational turn-taking, potentially leading to more realistic synthetic data.
@@ 2.3.3 On-Device Real-Time Inference
We implement our dual Llama models using the MLX framework Hannun et al. (2023) on the MPS device of Apple Silicon. The small model is quantized to bfloat16, while the large model is quantized to int8 for optimized performance. During streaming inference, both models continuously cache KV states for computational reuse. We use the testbench from MLX framework to measure the runtime and memory of dual Llama models in Apple M2 chip with 16GB memory. The token processed speed of the small model is 38.7 tokens/s and peak memory consumption is 2.49GB. The token generation speech of the large model is 14.2 tokens/s and the peak memory is 8.9GB. Our dual-model architecture achieves at least a 64% reduction in processing time during continuous inference.
@@ 3.1.1 Quantitative Metrics
To first measure whether LlamaPIE responds when it should, we compute its precision, recall, and accuracy (P/R/A) on our synthetic test set, using the data as ground truth for when assistance should be provided. In addition to hard P/R/A, we also report soft P/R/A, which gives model leniency of ±$\pm 1$ turn to respond, when the ground-truth data contains an assistance response. This is because responses can often be equally helpful when provided at many different nearby points in a conversation.
@@ 3.1.2 Qualitative Metrics
Due to the costly nature of large-scale human evaluation, we follow recent work and use LLM-as-a-judge Zheng et al. (2023). To improve the validity of our LLM evaluator, we use several key strategies. First, since we use Claude to generate our data, we used GPT-4o as our evaluator. Second, we apply a score rubric to our evaluation prompting, following suggestions from prior work Kim et al. (2023b).
@@
Validation with human judgment. To assess the quality of assistant responses, we evaluate each individual assistant response within a dialogue, comparing ratings between GPT-4o and human annotators. GPT-4o is instructed to assign a score to each response, while human annotators perform the same task based on a predefined rubric. Both LLM and human annotators classify responses into one of five categories: (1) not relevant/not used, (2) relevant but redundant/not needed, (3) relevant but not acted on, (4) relevant but used later, and (5) highly relevant/immediately used. These categories are worded to reduce ambiguity between how the evaluators rate responses. The full description of this rubric is shown in  §C.4.1. To mitigate hallucination and overly optimistic scoring, GPT-4o is also required to explain its reasoning in terms of both relevance and timeliness.
@@
To measure agreement between human and LLM evaluations, we compute the Pearson correlation coefficient between human annotators for each sample and compare it to the correlation between GPT-4o and a randomly selected human rating. The Pearson coefficient measures the linear correlation between two datasets, making it suitable for our 1–5 scoring system, where we expect a strong correlation along a 45-degree line.
We find that human-LLM correlation is stronger with $r=0.652$ compared to human-human correlation with $r=0.636$. Human validation is conducted only for the rubric scores, not for all nine principles, due to the time-intensive nature of human annotations. Nonetheless, the overall score for each assistant response correlates strongly with LLM assessment.
Accessing quality of synthetic datasets. We assessed the quality of our synthetic Claude-generated datasets using the above human validated GPT-based evaluator. The synthetic datasets achieved high rubric scores of 4.77, 4.78 and 4.88 for Synthetic, SODA, and PerLTQA, respectively. Across these datasets, the distribution of rubric scores was 0.68% with a score of 1, 0.51% with a score of 2, 6.83% with a score of 3, 1.45% with a score of 4 and 90.51% with a score of 5.
@@ 3.2 Small Model Accuracy Evaluation
We evaluate the small model’s ability to anticipate when it should respond. Table 2 shows that across all three dialogue test sets, hard precision and recall exceed 70%, while hard accuracy surpasses 93%. Additionally, soft precision and recall are around 90%, with soft accuracy exceeding 97%. These results indicate that the fine-tuned small model effectively anticipates when it should respond and that allowing a ±1 turn flexibility in response timing improves recall and precision.
@@ 3.3 Dual-Model Evaluation
Table 4: Ablation Study on MIT interview dataset. We evaluate on 100 conversation samples from the MIT dataset, consisting of a total of 1,324 turns. ‘Triggered by small’ represents the number of instances where the small model attempts to trigger assistance. ‘Responded by large’ represents the number of instances where the larger model provides actual assistance. Note that the ‘Triggered by small’ metric for the assistance-aware model is influenced by the large model, as the generated assistance is streamed back to the small model.
@@
Interestingly, the response frequency is significantly lower at approximately 5.8%. Unlike our synthetic datasets, which are designed with proactive assistance in mind, the MIT interviews involve natural conversations where users do not expect an in-ear assistant. This demonstrates that the proactive assistant remains silent when assistance is not anticipated, adapting appropriately to real-world scenarios.
3.3.1 Ablation study
We conduct ablation studies across different small and large model configurations on the MIT interview dataset. For the small model, we have two configurations: one finetuned to be aware of prior assistance provided by the assistant and the other is not aware of prior assistance. For the large model, we compare five different configurations: (1) Llama3.1-8B finetuned with no negative samples, (2) Llama3.1-8B finetuned with 25% negative samples, (3) Llama3.1-8B finetuned with 50% negative samples, (4) Llama3.1-1B finetuned with 25% negative samples, and (5) Llama3.1-8B with prompting and no fine-tuning. As shown in Table 4, if the small model is not aware of prior assistance, it will trigger assistance 2-3$\times$ more often. Moreover, being history-aware improves the overall response quality (rubric score). When we compare different large model configurations, we found that simply prompting the 8b model or finetuning the 1b model are both worse than finetuning 8b, showing that both the larger model size and finetuning are useful ingredients. Finally, the 8b model finetuned with 25% negative samples gave the highest rubric score, and we use this as our final model.
@@ 4 User Study: LlamaPIE in Real-Time Human Conversations
Each participant experiences three different conditions: Control, where no assistance is provided; Proactive (ours), where users receive assistance from LlamaPIE through the headset; Reactive (baseline), where users can access GPT-4o via a web UI during the conversation. In addition, five participants also experienced the following condition: Reactive (short audio), where users access GPT-4o via its voice mode during the conversation. GPT-4o was prompted to respond concisely (1-3 words) and users received assistance through the headset.
For each participant, we randomly assign a different topic to each of the three conditions. All participants begin with the control condition, while the order of the baseline LLM and proactive assistance conditions is randomized. Participants are given 3-5 minutes to read and memorize the background information for their assigned topic. The background information falls into two categories: (1) Wiki-style descriptions of information-dense topics such as quantum mechanics and DNA computing, and (2) Detailed profiles of fictional individuals, including their careers, families, hobbies, and personal interests. Further details are in §C.5.
@@ 4.1 Results
We collect a total of 50 dialogues. The average number of turns per dialogue is 11.2±${}_{\pm 1.97\,\text{(std)}}$. The average number of words per assistant response is 1.83, and the average response frequency is 25%.
@@
On the rubric score from  §C.4.1, participants give an average rating of 4.31±1.03. To gain deeper insights into response frequency, for Fig. 3(b), we ask participants to rate it using an opinion scale from -2 to +2:
  • -2: $>1$ assistance was unnecessary
  • -1: One assistance instance was unnecessary
@@
  • +2: $>1$ additional assistance was needed
@@
Overall, we find that human performance with LlamaPIE is on par to using a state-of-the-art LLM assistant, while overwhelmingly preserving the natural flow of the conversation as measured by both response latency and perceived disruption.
@@ 7 Limitations and Risks
Limitations. Currently, we provide memory only in the form of text for the user profile and prior conversations and events. Further research is needed to develop a system that automatically manages memory based on speech inputs (and potentially other sources). The models in our real-world prototype are trained solely on synthetic datasets; performance could likely be improved with real-world datasets, which are currently lacking. Our real-time prototype could eventually help to generate such datasets with human annotations. We have integrated open-source streaming ASR and TTS models into our dual-model pipeline to create a speech-to-speech setup. Instead of using cascaded systems, future work could integrate multi-modal speech models that directly process speech tokens to reduce latency. Additionally, incorporating personalization and human feedback could further enhance the customization of the in-ear assistant.

### v2

@@ Abstract
We introduce LlamaPIE, the first real-time proactive assistant designed to enhance human conversations through discreet, concise guidance delivered via hearable devices. Unlike traditional language models that require explicit user invocation, this assistant operates in the background, anticipating user needs without interrupting conversations. We address several challenges, including determining when to respond, crafting concise responses that enhance conversations, leveraging knowledge of the user for context-aware assistance, and real-time, on-device processing. To achieve this, we construct a semi-synthetic dialogue dataset and propose a two-model pipeline: a small model that decides when to respond and a larger model that generates the response. We evaluate our approach on real-world datasets, demonstrating its effectiveness in providing helpful, unobtrusive assistance. User studies with our assistant, implemented on Apple Silicon M2 hardware, show a strong preference for the proactive assistant over both a baseline with no assistance and a reactive AI assistant, highlighting the potential of LlamaPIE  to enhance live conversations. Code and dataset are available at https://github.com/chentuochao/LlamaPIE.
@@
Tuochao Chen  Nicholas Batchelder  Alisa Liu Noah A. Smith∗†Shyamnath Gollakota Paul G. Allen School of Computer Science & Engineering, University of Washington Allen Institute for Artificial Intelligence {tuochao,nicbat,alisaliu,nasmith,gshyam}@cs.washington.edu
@@ 1 Introduction
User interaction with language models (LMs) has primarily followed a turn-based dialogue format, where users actively request responses from LM assistants. This approach requires users to shift their attention and carefully phrase prompts to obtain useful answers. Thus, usability is restricted to scenarios where users can pause their activities to engage with the conversational assistant.
Figure 1: LlamaPIE is a Proactive, In-Ear assistant that augments human-to-human communication by providing discreet guidance via hearable devices. Its responses are short, provided only when helpful, and leverage the wearer’s memory of past events. In the figure, LlamaPIE assists Yuan by whispering 1–3 words to her during the conversation only when needed by anticipating user needs, remaining silent most of the time.
@@
Such proactive companions may enhance human interactions in a variety of contexts, including negotiations Deng et al. (2024), interviews Naim et al. (2018), customer support Deng et al. (2024), and cross-cultural communication Andolina et al. (2018). They may support neurodiverse individuals, such as those with autism or social anxiety, by assisting in the interpretation of social cues Liu et al. (2024), or serve as personal memory assistants for older adults experiencing progressive memory loss, providing contextually relevant information during conversations Bermejo et al. (2020).
Creating proactive assistants presents several challenges. Modern language models are typically designed to provide detailed responses, but in our case, assistance should be limited to just 1–3 words and offered only when it enhances the conversation without disrupting its flow. The assistant must also be context-aware, leveraging the wearer’s memory of prior activities and interactions. Achieving this enhancement to live conversations requires real-time streaming processing, while privacy considerations necessitate processing conversational data on edge devices rather than sending it to the cloud. Finally, there does not exist large-scale annotated data of in situ human conversations with real-time assistance for model training.
@@
Our work demonstrates a vision for proactive LM assistance that centers human-to-human conversation, and highlights the potential for such interactions to serve users in ways that complement the currently widespread human-AI “chat” interface.
@@ 2.2 Synthetic Dataset Generation
We construct examples by first creating the user memory (§2.2.2) and then the dialogue (§2.2.3). To improve the diversity and realism of the data, we draw from real conversational contexts and user profile datasets when synthetically generating data. The assistant should serve two main functions: providing reminders, e.g., helping users recall secondary details of events like names and places, and social guidance, e.g., helping the user continue the natural flow of the conversation. We use claude-3-5-sonnet-20240620 with Anthropic API to generate data (see §D.3 for dataset generation prompts).
@@ 2.2.2 Memory Generation or Selection
For synthetic memory, we use Claude to generate a user profile by providing five randomly selected keywords from a predefined list of 100 (see §D.2). Given these keywords, the model constructs a user profile and corresponding events. For SODA-based memory, instead of keywords, the model uses the context preceding SODA dialogues to generate the user profile. For PerLTQA-based memory, we select a random profile from the PerLTQA dataset, and two random events attached to that profile.
@@ 2.2.3 Dialogue Generation
The next step is dialogue generation, grounded in the user memory. We instruct Claude to construct a scenario that the user might encounter given their background. Each scenario is designed to fit a specified scenario type and illustrate a predefined use case, while also exemplifying two randomly selected principles of proactive assistants. To account for scenarios where users ignore assistance, some dialogues include an explicit instruction for Claude to generate interactions in which the user occasionally ignores the assistant’s messages (§D.3).
@@
Formatting for streaming and timing. To enable tokenized streaming and preserve timing information in dialogues — details typically absent in natural language data used for training LMs — we reformat Claude-generated dialogues to accommodate real-time processing. Gaps between speakers are replaced with silence markers, |SILENCE >, with each token representing 0.5 seconds of silence. Generated hesitation markers are replaced with silence markers as well (the hesitation duration is rounded up to the nearest equivalent number of tokens). This ensures that pauses and hesitations are accurately represented in the dataset, even when Claude’s generation does not include sufficiently long hesitations. Furthermore, since all gaps and hesitations are converted into the same silence marker, the streaming models do not have explicit marked information about hesitations. Future work could focus on more sophisticated modeling of conversational turn-taking, potentially leading to more realistic synthetic data.
@@ 2.3.3 On-Device Real-Time Inference
We implement our dual Llama models using the MLX framework Hannun et al. (2023) on the MPS device of Apple Silicon. The small model is quantized to bfloat16, while the large model is quantized to int8 for optimized performance. During streaming inference, both models continuously cache KV states for computational reuse. We use the testbench from MLX framework to measure the runtime and memory of dual Llama models in Apple M2 chip with 16GB memory. The token processed speed of the small model is 38.7 tokens/s and peak memory consumption is 2.49GB. The token generation speech of the large model is 14.2 tokens/s and the peak memory is 8.9GB. Our dual-model architecture achieves at least a 64% reduction in processing time during continuous inference compared to the implementation of a single large model.
@@ 3.1.1 Quantitative Metrics
To first measure whether LlamaPIE responds when it should, we compute its precision, recall, and accuracy (P/R/A) on our synthetic test set, using the data as ground truth for when assistance should be provided. In addition to hard P/R/A, we also report soft P/R/A, which gives model leniency of ±1$\pm$ 1± 1 turn to respond, when the ground-truth data contains an assistance response. This is because responses can often be equally helpful when provided at many different nearby points in a conversation.
@@ 3.1.2 Qualitative Metrics
Due to the costly nature of large-scale human evaluation, we follow recent work and use LM-as-a-judge Zheng et al. (2023). To improve the validity of our LM evaluator, we use several key strategies. First, since we use Claude to generate our data, we used GPT-4o as our evaluator. Second, we apply a score rubric to our evaluation prompting, following suggestions from prior work Kim et al. (2023b).
@@
Validation with human judgment. To assess the quality of assistant responses, we evaluate each individual assistant response within a dialogue, comparing ratings between GPT-4o and human annotators. GPT-4o is instructed to assign a score to each response, while human annotators perform the same task based on a predefined rubric. Both LM and human annotators classify responses into one of five categories: (1) not relevant/not used, (2) relevant but redundant/not needed, (3) relevant but not acted on, (4) relevant but used later, and (5) highly relevant/immediately used. These categories are worded to reduce ambiguity between how the evaluators rate responses. The full description of this rubric is shown in  §D.4.1. To mitigate hallucination and overly optimistic scoring, GPT-4o is also required to explain its reasoning in terms of both relevance and timeliness.
@@
To measure agreement between human and LLM evaluations, we compute the Pearson correlation coefficient between human annotators for each sample and compare it to the correlation between GPT-4o and a randomly selected human rating. The Pearson coefficient measures the linear correlation between two datasets, making it suitable for our 1–5 scoring system, where we expect a strong correlation.
We find that human-LM correlation is stronger with $r=0.652$ compared to human-human correlation with $r=0.636$. Human validation is conducted only for the rubric scores, not for all nine principles, due to the time-intensive nature of human annotations. Nonetheless, the overall score for each assistant response correlates strongly with LM assessment.
Assessing quality of synthetic datasets. We assessed the quality of our synthetic Claude-generated datasets using the above human validated GPT-based evaluator. The synthetic datasets achieved high rubric scores of 4.77, 4.78 and 4.88 for Synthetic, SODA, and PerLTQA, respectively. Across these datasets, the distribution of rubric scores was 0.68% with a score of 1, 0.51% with a score of 2, 6.83% with a score of 3, 1.45% with a score of 4 and 90.51% with a score of 5.
@@ 3.2 Small Model Accuracy Evaluation
We evaluate the small model’s ability to anticipate when it should respond. Table 2 shows that across all three dialogue test sets, hard precision and recall exceed 70%, while hard accuracy surpasses 93%. Additionally, soft precision and recall are around 90%, with soft accuracy exceeding 97%. These results indicate that the finetuned small model effectively anticipates when it should respond and that allowing a ±1 turn flexibility in response timing improves recall and precision.
@@ 3.3 Dual-Model Evaluation
Table 4: Ablation Study on MIT interview dataset. We evaluate on 100 conversation samples from the MIT dataset, consisting of a total of 1,324 turns. ‘Triggered by small’ represents the number of instances where the small model attempts to trigger assistance. ‘Responded by large’ represents the number of instances where the larger model provides actual assistance. Note that the ‘Triggered by small’ metric for the assistance-aware model is influenced by the large model, as the generated assistance is streamed back to the small model. The row with * corresponds to the results of the MIT interview dataset in Table.3
@@
Interestingly, the response frequency is significantly lower at approximately 5.8%. Unlike our synthetic datasets, which are designed with proactive assistance in mind, the MIT interviews involve natural conversations where users do not expect an in-ear assistant. This demonstrates that the proactive assistant often remains silent when assistance is not anticipated, adapting appropriately to real-world scenarios.
3.4 Ablation Study
We conduct ablation studies across different small and large model configurations on the MIT interview dataset. For the small model, we have two configurations: one finetuned to be aware of prior assistance provided by the assistant and the other is not aware of prior assistance. For the large model, we compare five different configurations: (1) Llama3.1-8B finetuned with no negative samples, (2) Llama3.1-8B finetuned with 25% negative samples, (3) Llama3.1-8B finetuned with 50% negative samples, (4) Llama3.1-1B finetuned with 25% negative samples, and (5) Llama3.1-8B with prompting and no fine-tuning. As shown in Table 4, if the small model is not aware of prior assistance, it will trigger assistance 2–3$\times$ more often. Moreover, being history-aware improves the overall response quality (rubric score). When we compare different large model configurations, we found that simply prompting the 8B model or finetuning the 1B model are both worse than finetuning 8B, showing that both the larger model size and finetuning are useful ingredients. Finally, the 8B model finetuned with 25% negative samples gave the highest rubric score, and we use this as our final model.
@@ 4 User Study: LlamaPIE in Real-Time Human Conversations
Each participant experiences three different conditions: Control, where no assistance is provided; Proactive (ours), where users receive assistance from LlamaPIE through the headset; Reactive (baseline), where users can access GPT-4o via a web UI during the conversation. In addition, five participants also experienced the following condition: Reactive (short audio), where users access GPT-4o via its voice mode during the conversation. GPT-4o was prompted to respond concisely (1–3 words) and users received assistance through the headset.
For each participant, we randomly assign a different topic to each of the three conditions. All participants begin with the control condition, while the order of the baseline LM and proactive assistance conditions is randomized. Participants are given 3–5 minutes to read and memorize the background information for their assigned topic. The background information falls into two categories: (1) Wiki-style descriptions of information-dense topics such as quantum mechanics and DNA computing, and (2) Detailed profiles of fictional individuals, including their careers, families, hobbies, and personal interests. Further details are in §D.5.
@@ 4.1 Results
[…]
@@
On the rubric score from §D.4.1, participants give an average rating of 4.31±1.03. To gain deeper insights into response frequency, for Fig. 3(b), we ask participants to rate it using an opinion scale from -2 to +2:
  • –2: >1>1> 1 assistance was unnecessary
  • –1: One assistance instance was unnecessary
@@
  • +2: >1>1> 1 additional assistance was needed
@@
Overall, we find that human performance with LlamaPIE is on par to using a state-of-the-art LM assistant, while overwhelmingly preserving the natural flow of the conversation as measured by both response latency and perceived disruption.
@@ 7 Limitations and Risks
Limitations. Currently, we provide memory only in the form of text for the user profile and prior conversations and events. Further research is needed to develop a system that automatically manages memory based on speech inputs (and potentially other sources). The models in our real-world prototype are trained solely on synthetic datasets; performance could likely be improved with real-world datasets, which are currently lacking. Our real-time prototype could eventually help to generate such datasets with human annotations. We have integrated open-source streaming ASR and TTS models into our dual-model pipeline to create a speech-to-speech setup. Instead of using cascaded systems, future work could integrate multimodal speech models that directly process speech tokens to reduce latency. Additionally, incorporating personalization and human feedback could further enhance the customization of the in-ear assistant.